# Learning to flock in open space by avoiding collisions and staying together

### v1

@@ Appendix A Flock Structure
with the adjacency matrix $W_{ij}(t)=1$ if agents $i$ and $j$ are topological neighbours at time $t$ and 0 otherwise. A non-zero mixing rate implies neighbour exchange and an irreversible, nonequilibrium dynamics [6]. Indeed, it can be shown that the entropy production rate of Vicsek-like models is proportional to such a mixing rate [32]. The mixing rates measured in our trained flocks are small but finite, $\mu\approx 5\cdot 10^{-4}$ for the typical parameters considered in this work. Particles do flow, but their neighbour rearrangement occurs on a much slower timescale than local alignment, consistently with the state of local quasi-equilibrium observed in starling flocks [23].
@@
A typical pair distribution function for a trained flock of $N=400$ agents is reported in Fig. 5a. It presents a single peak at $r\approx 0.5$ and no further structure, in qualitative agreement with experimental observation in starling flocks [17, 24] whose $g(r)$ also lacks the oscillating behavior of typical liquids. The only peak we observe is a consequence of a well defined peak in the mean neighbour distance distribution (see Fig. 5b) and is due to the effective repulsion discussed in Sec. 3.2 and a further repulsive effect due to noisy fluctuations in the orientation [36].
Appendix B Policy distillation
Figure 6: Representation of the average $Q$-matrix $\bar{Q}=\langle Q_{i}\rangle_{i}$ trained in the DT scheme for $N=100$ agents. Each plot, labeled by an integer $n$ refers to a specific state $d=n\Delta d$ and shows $\bar{Q}(d,\beta)$ as a function of $\beta$. Error bars represent one standard error.
A detailed analysis of the trained (and averaged over all agents for the DT scheme) $Q$-matrix state-by-state (see for instance Fig. 6) shows that they are characterized by a well-defined minimum only for small mean neighbour distances. This shows that for $d<d^{*}$ training strongly converges towards a well-defined optimal policy of fully aligning interactions (with the only exception of the case $d<v_{0}$ as discussed in Sec. 3.2). For $d>d^{*}$, on the other hand, absolute minima are not well defined, and indeed one can verify that at large distances different training instances do not always converge towards the same optimal policy $\beta^{*}(d)$. This suggests that, as long as some degree of attractive interaction is present, the exact value of the policy $\beta^{*}(d)$ is unimportant at large distances.
@@ Appendix B Policy distillation
An analogous policy distillation can be performed to remove strong attraction arising for $d<v_{0}$, as shown in the top left panel of Fig. 6. We have verified that substituting the attractive optimal policy selected by training with pure alignment ($\beta=0$) still results in cohesive flocking (see Fig. 7, red curves), albeit with a slightly larger order parameter and smaller flock size due to the lack of the effective repulsion discussed in Sec. 3.2).
Figure 7: Time-series of the order parameter $\Phi$ (a) and flock size $\Gamma$ for $N=100$ agents with two different policy distillations (see text): Random policy for $d>d^{*}$ (blue curves) and complete aligning interaction for $d<v_{0}$ (red curves).

### v2

@@ III.2 Flock Structure
with the adjacency matrix $W_{ij}(t)=1$ if agents $i$ and $j$ are topological neighbours at time $t$ and 0 otherwise. A non-zero mixing rate implies neighbour exchange and an irreversible, nonequilibrium dynamics [21]. Indeed, it can be shown that the entropy production rate of Vicsek-like models is proportional to such a mixing rate [18]. The mixing rates measured in our trained flocks are small but finite, $\mu\approx 5\cdot 10^{-2}$ for the typical parameters considered in this work. Particles do flow, but their neighbour rearrangement occurs on a much slower timescale than local alignment, consistently with the state of local quasi-equilibrium observed in starling flocks [30].
@@
A typical pair distribution function for a trained flock of $N=400$ agents is reported in Fig. 3a. It presents a single peak at $r\approx 0.5$ and no further structure, in qualitative agreement with experimental observation in starling flocks [9, 8] whose $g(r)$ also lacks the oscillating behavior of typical liquids. The only peak we observe is a consequence of a well defined peak in the mean neighbour distance distribution (see Fig. 3b) and is due to the effective repulsion discussed in Sec. III.3 and a further repulsive effect due to noisy fluctuations in the orientation [6].
III.3 Analysis of the Q-matrix
The (normalized) final $Q$-matrices for CT and DT are shown (respectively) in Fig. 4a and Fig. 4c for $N=100$. They are characterized by two main regimes, a first one for states $d>d^{*}\approx 2$ and a second one at shorter distances, $d<d^{*}$. At large mean distances to neighbours, the optimal actions for the agents, given by Eq. (7), involve a combination of alignment with, and attraction to, their neighbours, with $\beta$ seemingly randomly scattered in the interval $(0,1]$.
Note finally the exception of the first state, $d<\Delta d=0.2$, where attraction seems to be the optimal action in both CT and DT. This rather puzzling result is a consequence of the discrete nature of our dynamics: when $d<\bar{d}$ the cost function rewards actions that increase the mean distance to neighbours $d$ and penalizes those that further reduce it. For $d>v_{0}$ this simply rules out attraction, resulting in agents with commonly aligned headings that separate slowly due to noisy fluctuations in their individual headings. However, when extremely close, $d\leq v_{0}$, strong attraction may lead to large differences in neighbours’ heading, so that nearby agents may easily pass each other and find themselves, after a single time-step, separated by a larger distance. In practice, at distances smaller than $v_{0}$, large $\beta$ may act like an effective repulsion and be favored by the diverging cost function. See also Section III.D for further evidence
We have indeed verified that, for a reduced speed $v_{0}\ll\Delta d$ (in particular, we have chosen $v_{0}=0.05$), this effect is largely mitigated, with a typical optimal policy for the smallest mean neighbour distance state ($d<\Delta d$) dominated by alignment, $\beta\in[0,0.2]$. In Section III.4 we further show that forcing $\beta=0$ for $d<v_{0}$ still leads to cohesive flocking.
III.4 Policy distillation
Figure 5: Representation of the average $Q$-matrix $\bar{Q}=\langle Q_{i}\rangle_{i}$ trained in the DT scheme for $N=100$ agents. Each curve, labeled by an integer $n$ refers to a specific state $d=n\Delta d$ and shows $\bar{Q}(d,\beta)$ as a function of $\beta$. Error bars represent one standard error. (a) First state, $d<v0$. (b) States with $d<d^{*}$. (c) States with $d>d^{*}$. Note that the ratio between the vertical scale of the three panels is 1:2:40.
A detailed analysis of the trained (and averaged over all agents for the DT scheme) $Q$-matrix state-by-state (see for instance Fig. 5) shows that they are characterized by a well-defined minimum only for small mean neighbour distances. This shows that for $d<d^{*}$ training strongly converges towards a well-defined optimal policy of fully aligning interactions (with the only exception of the case $d<v_{0}$ as discussed in Sec. III.3). For $d>d^{*}$, on the other hand, absolute minima are not well defined, and indeed one can verify that at large distances different training instances do not always converge towards the same optimal policy $\beta^{*}(d)$. This suggests that, as long as some degree of attractive interaction is present, the exact value of the policy $\beta^{*}(d)$ is unimportant at large distances.
@@ III.4 Policy distillation
An analogous policy distillation can be performed to remove strong attraction arising for $d<v_{0}$, as shown in the top left panel of Fig. 5. We have verified that substituting the attractive optimal policy selected by training with pure alignment ($\beta=0$) still results in cohesive flocking (see Fig. 6, red curves), albeit with a slightly larger order parameter and smaller flock size due to the lack of the effective repulsion discussed in Sec. III.3.
Figure 6: Time-series of the order parameter $\Phi$ (a) and flock size $\Gamma$ (b) for $N=100$ agents with two different policy distillations (see text): Random policy for $d>d^{*}$ (blue curves) and complete aligning interaction for $d<v_{0}$ (red curves).
III.5 A swarming counterexample
We conclude by discussing the role of the cost function (8) divergence at short distances in the onset of flocking behavior. Such a divergence can be removed by considering instead $\mathcal{L}(d+z)$ with a positive shift $z$. In particular, with the choice $z=\bar{d}=1$ the minimum of the cost function is shifted to $d=0$, completely removing penalties for arbitrarily small agent distances $d$. We have verified, as shown in Fig. 7, that this choice suppresses flocking and only leads to cohesive swarming (see Video2 in Supplementary). The analysis of the trained $Q$-matrix indeed shows that in this case the orientational dynamics is dominated by attraction ($\beta\approx 1$).
III.6 Boundary agents and Voronoi based cost
Since Voronoi cells at the group boundary have, by definition, an infinite area, corresponding to a vanishing cost function $\mathcal{L}$, our new choice highlights the cost/reward difference between agents located inside and those at the boundary of the flock. Therefore, actions resulting in agents moving from the boundary to the interior of the group will be strongly rewarded, while moving from the interior to the border is strongly penalized. Moreover, agents at the boundary get zero cost/reward for any action that does not move them to the inside of the flock, regardless of any change in the distance to their Voronoi neighbours. We have numerically tested this new learning scheme both for centralized and decentralized training, but we have not found any qualitative difference with cost/rewards based on the distance to neighbours, with agents still learning similar optimal policies leading to cohesive flocking, as shown for instance in Fig. 8.
Figure 8: Comparison between the optimal policy dynamics of flocks of $N=100$ agents trained in the CT paradigm with cost functions based on mean distance (blue) or Voronoi area (red). (a) Scalar order parameter vs. time. (b) Flock size vs. time.
IV Discussion
Interestingly, the learned flocking dynamics shows characteristic structural features, such as a slow mixing rate and a structureless pair distribution function, reminiscent of natural bird flocks, reinforcing the biological plausibility of the learned behavior. While our results have been obtained in two spatial dimensions, one should remember that, due to the presence of gravity, real birds fly preferentially in the horizontal directions, resulting in a kind of quasi-2d motion [2, 36]. For this reason, we believe our results may be of direct relevance for real birds.
We finally note that our results may also prove useful beyond animal behavior, particularly in the field of swarm robotics [24]. In the future, it may be interesting to generalize these results to mixed-agent populations and environments with external threats/perturbations in order to extend the result of [12] to the full non-equilibrium dynamics learned by our agents.
V Acknowledgments